OpenStudy (anonymous):

DIFFERENTIAL EQUATION: help please, anybody. (D^2 +4)y =4sin^2 (x)

OpenStudy (shubhamsrg):

y= a cos2x ?

OpenStudy (shubhamsrg):

guess that substn should work..

OpenStudy (anonymous):

aaa. i think i got it. wait, i'll try.

OpenStudy (shubhamsrg):

ohh wait,,i didnt see a "y" also in there. !!

OpenStudy (anonymous):

@shubhamsrg I think it's just written in 'operator' notation... \[\frac{ d^2y }{dt^2 } +4y = 4\sin^2x\] correct @lilMissMindset ?

OpenStudy (anonymous):

heh.

OpenStudy (shubhamsrg):

there's a y also with D^2 .. i had got that,,i didnt see a y being multiplied to whole that time..sorry..

OpenStudy (anonymous):

\[\frac{ d^2y }{ dx^2 } +4y =4\sin^2x\] rather.

OpenStudy (anonymous):

yes, tat's it

OpenStudy (anonymous):

that*

OpenStudy (anonymous):

yes he's correct I believe y =a*cos^2 x

OpenStudy (anonymous):

did you get it to work?

OpenStudy (anonymous):

that's the particular soln. anyway...

OpenStudy (anonymous):

isn't it like, Acosx+Bsinx+Cxcosx+Dxsinx

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

since sinx is twice? applying case 4 of undetermined coefficients

OpenStudy (anonymous):

where it has some repeated imaginary root

OpenStudy (anonymous):

look d^2 (a*cos^2(x)) /dx^2 +4a cos^2(x) = 4sin^2(x) second derivative of a*cos^2(x) is 2a*cos^2(x) - 2a*sin^2(x)

OpenStudy (anonymous):

this is for the particular solution...

OpenStudy (anonymous):

finding the homogeneous solution shouldn't be a problem...

OpenStudy (anonymous):

\[\lambda ^2 +4 =0\]

OpenStudy (anonymous):

lambda = 2i

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

i'll work on it. please guide me still, in case i get stuck.

OpenStudy (anonymous):

hold on... can you tell me what part is giving you problems?

OpenStudy (anonymous):

getting Yp gives me a lot of trouble. since you told me that it's acos^2 x, i am differentiating it now.

OpenStudy (anonymous):

ok:)

OpenStudy (anonymous):

how about if the given is cos^x, then Yp=asin^2 x?

OpenStudy (anonymous):

if the forcing term is cos^2 x you mean?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

i forgot 2, lol.

OpenStudy (anonymous):

yeah, for the same equation, I believe that a*sin^2 x would be the trial solution then...

OpenStudy (anonymous):

yes. it would. just checked it.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

sure:)

OpenStudy (anonymous):

darn, i can't do it. show me a brief solution please. so i can learn how.

OpenStudy (anonymous):

just for the Yp part. swear.

OpenStudy (anonymous):

well, I did it already (a*cos^2(x))" +4a cos^2(x) = 4sin^2(x)

OpenStudy (anonymous):

2a*cos^2(x) - 2a*sin^2(x) +4a cos^2(x) = 4sin^2(x)

OpenStudy (anonymous):

hmm, I see your point ...

OpenStudy (anonymous):

I guess you'll be needing something like yp = a*cos^2(x) + b*sin^2(x)

OpenStudy (anonymous):

let me try it out and see...

OpenStudy (anonymous):

sure. :)

OpenStudy (anonymous):

seems to work... I got a= 2 b=-2

OpenStudy (anonymous):

so 2cos^2 x -2sin^2 x

OpenStudy (anonymous):

answer here in the book is 1/2(1-xsin2x) for the Yp part.

OpenStudy (anonymous):

I'll assume that's (1/2)*( 1 - x*sin^2 x )

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

you should check the original, I think you might have written the problem wrong...

OpenStudy (anonymous):

you left out and x I think...

OpenStudy (anonymous):

i don't think so.

OpenStudy (anonymous):

(1/2)*( sin^2 x - x*sin 2x ) ??

OpenStudy (anonymous):

I don't see how (1/2)(1-xsin2x) can work... maybe I'm missing something.

OpenStudy (anonymous):

something's off in your original equation or that soln.